if a+b+C=0and magnitude of a=b and magnitude of c is root 2 times of a and b then angle between the  vectors

Grade:11

1 Answers

Vikram
5 years ago
it clearly indicates that A&B represents adjacent sides of a square with C as it’s diagonal.
now angle between A&B is 90. So angle between A&C  or B&C is 270/2= 135 degree.
It is given that  a+b+c=0. That means a+b=_c. Also given that a=b and c is their resultant vector having magnitude root 2times.